# Optimally truncated WKB approximation for the 1D stationary Schrödinger equation in the highly oscillatory regime

### v1

@@ Abstract
This paper is dedicated to the efficient numerical computation of solutions to the 1D stationary Schrödinger equation in the highly oscillatory regime. We compute an approximate solution based on the well-known WKB-ansatz, which relies on an asymptotic expansion w.r.t. the small parameter $\varepsilon$. Assuming that the coefficient in the equation is analytic, we derive an explicit error estimate for the truncated WKB series, in terms of $\varepsilon$ and the truncation order $N$. For any fixed $\varepsilon$, this allows to determine the optimal truncation order $N_{opt}$ which turns out to be proportional to $\varepsilon^{-1}$. When chosen this way, the resulting error of the optimally truncated WKB series behaves like $\mathcal{O}(\varepsilon^{-2}\exp(-r/\varepsilon))$, with some parameter $r>0$. The theoretical results established in this paper are confirmed by several numerical examples.
@@ 1 Introduction
In this paper we are concerned with the numerical solution of the highly oscillatory and stationary 1D Schrödinger equation
$\begin{cases}\varepsilon^{2}\varphi^{\prime\prime}(x)+a(x)\varphi(x)=0\>,\quad x\in I:=[\xi,\eta]\>,\\ \varphi(\xi)=\varphi_{0}\>,\\ \varepsilon\varphi^{\prime}(\xi)=\varphi_{1}\>.\end{cases}$ (1)
Here, $0<\varepsilon\ll 1$ is a very small parameter and $a$ is a real-valued function satisfying $a(x)\geq a_{0}>0$ and, for a quantum mechanical problem, it is related to the potential. The constants $\varphi_{0},\varphi_{1}\in\mathbb{C}$ may depend on $\varepsilon$ but are assumed to be ${\varepsilon}$-uniformly bounded. It is known that the (local) wave length $\lambda$ of the solution $\varphi$ to (1) is proportional to $\varepsilon$. More precisely, it can be expressed as $\lambda(x)=(2\pi\varepsilon)/\sqrt{a(x)}$. Consequently, for a small parameter $\varepsilon$ the solution becomes highly oscillatory, particularly in the semi-classical limit $\varepsilon\to 0$.
Highly oscillatory problems such as (1) occur across a broad range of applications, e.g., plasma physics [10, 15], inflationary cosmology [17, 25] and electron transport in semiconductor devices such as resonant tunneling diodes [19, 23, 20]. More specifically, the state of an electron of mass $m$ that is injected with the prescribed energy $E$ from the right boundary into an electronic device (e.g., diode), modeled on the interval $[\xi,\eta]$, can be described by the following boundary value problem (BVP) (e.g., see [3] or [20, Chap. 2]):
$\begin{cases}-\varepsilon^{2}\psi_{E}^{\prime\prime}(x)+V(x)\psi_{E}(x)=E\psi_{E}(x)\>,\quad x\in(\xi,\eta)\>,\\ \psi_{E}^{\prime}(\xi)+\operatorname{i}k(\xi)\psi(\xi)=0\>,\\ \psi_{E}^{\prime}(\eta)-\operatorname{i}k(\eta)\psi(\eta)=-2\operatorname{i}k(\eta)\>.\end{cases}$ (2)
@@ 1.1 Background and approach
Since the solution $\varphi$ to (1) exhibits rapid oscillations when $\varepsilon$ is small, standard numerical methods for ODEs become inefficient as they are typically constrained by grid limitations $h=\mathcal{O}(\varepsilon)$ ($h$ denoting the step size), in order to resolve the oscillations accurately. By contrast, in [16, 13] uniformly accurate (w.r.t. $\varepsilon$) marching methods were proposed which yield global errors of order $\mathcal{O}(h^{2})$ and allow to reduce the grid limitation to at least $h=\mathcal{O}(\sqrt{\varepsilon})$. The WKB-based (named after the physicists Wentzel, Kramers, Brillouin; cf. [14]) one-step method from [3] is even asymptotically correct, i.e. the numerical error goes to zero with $\varepsilon\to 0$, provided that the integrals $\int^{x}\sqrt{a(\tau)}\,\mathrm{d}\tau$ and $\int^{x}a(\tau)^{-1/4}(a(\tau)^{-1/4})^{\prime\prime}\,\mathrm{d}\tau$ for the phase of the solution can be computed exactly. More precisely, the method then yields an error which is of order $\mathcal{O}(\varepsilon^{3})$ as $\varepsilon\to 0$ and $\mathcal{O}(h^{2})$ as $h\to 0$. If these integrals cannot be evaluated exactly, the asymptotically correct error behavior can be (almost) recovered by employing spectral methods for the integrals, as shown in [4]. Further, in [2] the authors propose a numerical algorithm, which switches adaptively between a defect correction iteration (which builds on an asymptotic expansion) for oscillatory regions of the solution, and a conventional Chebyshev collocation solver for smoother regions. Although the method is demonstrated to be highly accurate and efficient, a full error analysis was left for future work.
@@
The key question when implementing this WKB approximation is which choice of $N$ is adequate or even optimal, in the sense of minimizing the resulting approximation error. Indeed, since the asymptotic WKB series is typically divergent, the error cannot simply be reduced further by increasing the value of $N$. This question about the best attainable accuracy of the WKB approximation was already addressed in [25], where the author compared the WKB series with the exact solution represented by a convergent Bremmer series, or more precisely, by an asymptotic expansion of that Bremmer series. The author finds that in cases where the coefficient function $a$ is analytic, the optimal truncation order is proportional to $\varepsilon^{-1}$, yielding a corresponding optimal accuracy which is exponentially small w.r.t. $\varepsilon$. However, to derive these results, the author makes several additional asymptotic approximations. In the present paper, on the other hand, we shall follow a more rigorous strategy by providing error estimates for the WKB approximation which are explicit w.r.t. $\varepsilon$ and $N$. We note, however, that the key assumption from [25], i.e., $a$ being analytic, will also be crucial for the strategy of the present work.
@@ 1.2 Main results
Our first main result is Theorem 3.7, which provides an explicit (w.r.t. $\varepsilon$ and $N$) error estimate for the WKB approximation, and implies that the approximation error is of order $\mathcal{O}(\varepsilon^{N})$. The explicitness of this estimate then allows the investigation of the error w.r.t. the truncation order $N$. Indeed, the optimal truncation order $N_{opt}$ can be predicted by minimizing the established upper error bound w.r.t. $N$ or by determining the smallest term of the asymptotic series, and is found to be proportional to $\varepsilon^{-1}$. This leads to our second main result, namely, Corollary 4.1. It states that, for an adequate choice of $N=N(\varepsilon)\sim\varepsilon^{-1}$, the error of the WKB approximation is of order $\mathcal{O}(\varepsilon^{-2}\exp(-r/\varepsilon))$, $r>0$ being some constant. As a consequence, also the error of the optimally truncated WKB approximation is of order $\mathcal{O}(\varepsilon^{-2}\exp(-r/\varepsilon))$, see also Remark 4.2.
@@ 2 WKB approximation
According to [5, Sec. 10.2], for the WKB-ansatz (5)-(6) to be valid on the whole interval $I$, it is necessary that the series $\sum_{n=0}^{\infty}\varepsilon^{n-1}S_{n}(x)$ is a uniform asymptotic expansion of $\varepsilon^{-1}S(x)$ as $\varepsilon\to 0$. This implies that for any $n\in\mathbb{N}_{0}$ the relation
@@ 3 Error analysis
with constants $\hat{\varphi}_{0},\hat{\varphi}_{1}\in\mathbb{C}$. Then there exists $C>0$ independent of $\varepsilon$ such that
@@
$\frac{\mathrm{d}}{\mathrm{d}x}E(x)$ $\leq\frac{1}{\varepsilon^{2}}|f|^{2}+cE(x)\>.$ (25)
Applying Gronwall’s inequality, we therefore get
$E(x)$ $\leq\left(\frac{1}{\varepsilon^{2}}\lVert f\rVert_{L^{2}(I)}^{2}+E(\xi)\right)\operatorname{e}^{c(x-\xi)}$
@@
In order to derive an error estimate for the WKB approximation (13), which is explicit not only w.r.t. $\varepsilon$ but also w.r.t. the truncation order $N$, it is essential to control the growth of the functions $S_{n}$ w.r.t. $n\in\mathbb{N}_{0}$. As a first step, we aim to establish upper bounds for the derivatives $S_{n}^{\prime}$ which are given by recurrence relation (8)-(10). To this end, we employ a strategy similar to [18, Lemma 2], which relies heavily on Cauchy’s integral formula. To enable us to apply this tool, we shall assume that $S_{0}^{\prime}$ is not only defined on the real interval $I$, but also on a complex neighbourhood $G\subset\mathbb{C}$ of $I$. This leads us to introduce the following assumption.
@@
Let $S_{0}^{\prime}$ be analytic on a complex, bounded, simply connected neighbourhood $G\subset\mathbb{C}$ of $I$, satisfying $S_{0}^{\prime}(z)\neq 0$ for any $z\in G$.
As a consequence of Hypothesis A, $a$ and all $S_{n}$, $n\in\mathbb{N}$ are analytic on $G$. In particular, each $S_{n}$ is bounded on $I$.
For the next lemma, we introduce the open set
$G_{\delta}:=\{z\in G\mid\operatorname{dist}(z,\partial G)>\delta\}$ (27)
for some $\delta>0$.
@@
Let Hypothesis A be satisfied and let $0<\delta\leq 1$ be such that $G_{\delta}\neq\emptyset$. Then there exist constants $0<\widehat{K}_{\delta}\leq\widehat{K}$ with $\widehat{K}$ depending only on $G$ and $S_{0}^{\prime}$ such that
$\lVert S_{n}^{\prime}\rVert_{L^{\infty}(G_{\delta})}\leq\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}\widehat{K}_{\delta}^{n}n^{n}\delta^{-n}\>,\quad n\in\mathbb{N}_{0}\>.$ (28)
Here we define $0^{0}$ as $1$.
@@
Note that since $S_{0}^{\prime}$ is analytic on $G$, it follows from recurrence relation (8)-(10) that $S_{n}^{\prime}$ and hence also $\widehat{S}_{n}^{\prime}$ are analytic on $G$, for every $n\in\mathbb{N}_{0}$. We will now prove by induction on $n$ that
$\lVert\widehat{S}_{n}^{\prime}\rVert_{L^{\infty}(G_{\delta})}\leq\frac{1}{2}\widehat{K}_{\delta}^{n}n^{n}\delta^{-n}\>,\quad n\in\mathbb{N}_{0}\>.$ (31)
Obviously, this estimate does hold for $n=0$, according to (29). Assume now that the estimate in (31) holds for $10\leq j\leq n-$ 1 with some fixed $n\geq 1$. We will now prove it for $n$. Let $10<\kappa<$ 1 and $z\in G_{\delta}$. We denote with $\partial B_{\kappa\delta}(z)$ a circle of radius $\kappa\delta$ around $z$, see the left of Figure 1.
@@
This, together with the fact that $\partial B_{\kappa\delta}(z)\subseteq G_{(1-\kappa)\delta}$, now yields
@@
$\leq\frac{1}{4}\widehat{K}_{\delta}^{n}\delta^{-n}\sum_{j=1}^{n-1}j^{j}(n-j)^{n-j}$
$\quad+\frac{1}{4}\lVert(S_{0}^{\prime})^{-2}\rVert_{L^{\infty}(G_{\delta})}\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}\delta^{-n}\widehat{K}_{\delta}^{n-1}\frac{(n-1)^{n-1}}{\kappa(1-\kappa)^{n-1}}\>.$ (34)
@@
$\lVert\widehat{S}_{n}^{\prime}\rVert_{L^{\infty}(G_{\delta})}$ $\leq\frac{1}{2}\widehat{K}_{\delta}^{n}n^{n}\delta^{-n}\left[\frac{1}{2}\left(\frac{n-1}{n}\right)^{n}+\frac{\lVert(S_{0}^{\prime})^{-2}\rVert_{L^{\infty}(G_{\delta})}\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}}{2\widehat{K}_{\delta}\kappa(1-\kappa)^{n-1}n}\left(\frac{n-1}{n}\right)^{n-1}\right]\>.$ (35)
@@
$\lVert\widehat{S}_{n}^{\prime}\rVert_{L^{\infty}(G_{\delta})}$ $\leq\frac{1}{2}\widehat{K}_{\delta}^{n}n^{n}\delta^{-n}\left[\frac{1}{2\operatorname{e}}+\frac{\lVert(S_{0}^{\prime})^{-2}\rVert_{L^{\infty}(G_{\delta})}\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}}{2\widehat{K}_{\delta}}\right]$ (36)
@@
$\widehat{K}_{\delta}:=\frac{\operatorname{e}}{2\operatorname{e}-1}\lVert(S_{0}^{\prime})^{-2}\rVert_{L^{\infty}(G_{\delta})}\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}\>.$ (37)
Finally, we note that $\widehat{K}_{\delta}\leq\widehat{K}:=\frac{\operatorname{e}}{2\operatorname{e}-1}\lVert(S_{0}^{\prime})^{-2}\rVert_{L^{\infty}(G)}\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}$ for any $0<\delta\leq 1$. This concludes the proof. ∎
@@
By applying Lemma 3.2 on the r.h.s. of (40) we conclude that (39) holds with $K_{1}:=1/\delta$ and $K_{2}:=\widehat{K}_{\delta}/\delta$. Estimate (38) then follows from (39) for $k=1$ and by the definition of $S_{n}$, see (14). ∎
@@
Of course, it is of great interest to find a constant $K_{2}$ from Corollary 3.3 which is as small as possible. To this end one would have to minimize the constant $\widehat{K}_{\delta}/\delta$ in estimate (28). In particular, one has to fix some complex neighbourhood $G$ of $I$ as well as a constant $0<\delta\leq 1$ such that it holds $I\subset G_{\delta}$. Further, the proof of Lemma 3.2 indicates that $\widehat{K}_{\delta}$ can be reduced by choosing $G$ small, see (37). However, this means that one is forced to reduce also the value of $\delta$. Hence, this procedure usually results in a trade-off between the magnitudes of $\widehat{K}_{\delta}$ and $\delta$. More precisely, one would have to solve the following minimization problem:
$\min_{\begin{subarray}{c}0<\delta\leq 1\\ G\subset\mathbb{C}\\ I\subset G_{\delta}\end{subarray}}\frac{\lVert(S_{0}^{\prime})^{-2}\rVert_{L^{\infty}(G_{\delta})}\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}}{\delta}=\lVert(S_{0}^{\prime})^{-2}\rVert_{L^{\infty}(I)}\min_{\begin{subarray}{c}0<\delta\leq 1\\ G^{\delta}\subset\mathbb{C}\end{subarray}}\frac{\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G^{\delta})}}{\delta}\>,$ (46)
where $G^{\delta}:=\{z\in\mathbb{C}\mid\operatorname{dist}(z,I)<\delta\}$. Equality in (46) holds for the following reasons: First, on the l.h.s. of (46) one only needs to consider sets $G\subset\mathbb{C}$ of the form $G=G^{\kappa}:=\{z\in\mathbb{C}\mid\operatorname{dist}(z,I)<\kappa\}$, with $\kappa>0$. For any $0<\delta\leq 1$ such that $I\subset G_{\delta}$, this follows since the numerator on the l.h.s. of (46) is not increased when replacing $G$ and $G_{\delta}$ by $G^{\operatorname{dist}(I,\partial G)}$ and $(G^{\operatorname{dist}(I,\partial G)})_{\delta}$, respectively, see the right of Figure 1. The condition $I\subset G_{\delta}$ then simply reads $\kappa>\delta$. Second, since for a fixed $0<\delta\leq 1$ and $\delta<\kappa_{1}<\kappa_{2}$ it holds that $G^{\kappa_{1}}\subset G^{\kappa_{2}}$ and $(G^{\kappa_{1}})_{\delta}\subset(G^{\kappa_{2}})_{\delta}$, it is sufficient to consider simply the sets $G^{\delta+\epsilon}$, with $\epsilon>0$ being an arbitrarily small number. The equality in (46) then follows from the fact that $\lim_{\epsilon\to 0}G^{\delta+\epsilon}=G^{\delta}$ and $\lim_{\epsilon\to 0}(G^{\delta+\epsilon})_{\delta}=\lim_{\epsilon\to 0}G^{\epsilon}=\bigcap_{\epsilon>0}G^{\epsilon}=I$.
@@
$=\widetilde{\varphi}_{N}\left(\sum_{0\leq n,k\leq N}\varepsilon^{n+k}S_{n}^{\prime}S_{k}^{\prime}+\sum_{n=0}^{N}\varepsilon^{n+1}S_{n}^{\prime\prime}+a(x)\right)\>.$ (49)

### v2

@@ Abstract
This paper is dedicated to the efficient numerical computation of solutions to the 1D stationary Schrödinger equation in the highly oscillatory regime. We compute an approximate solution based on the well-known WKB-ansatz, which relies on an asymptotic expansion w.r.t. the small parameter $\varepsilon$. Assuming that the coefficient in the equation is analytic, we derive an explicit error estimate for the truncated WKB series, in terms of $\varepsilon$ and the truncation order $N$. For any fixed $\varepsilon$, this allows to determine the optimal truncation order $N_{opt}$ which turns out to be proportional to $\varepsilon^{-1}$. When chosen this way, the resulting error of the optimally truncated WKB series behaves like $\mathcal{O}(\exp(-r/\varepsilon))$, with some parameter $r>0$. The theoretical results established in this paper are confirmed by several numerical examples.
@@ 1 Introduction
In this paper we are concerned with the numerical solution of the stationary 1D Schrödinger equation
$\begin{cases}\varepsilon^{2}\varphi^{\prime\prime}(x)+a(x)\varphi(x)=0\>,\quad x\in I:=[\xi,\eta]\>,\\ \varphi(\xi)=\varphi_{0}\>,\\ \varepsilon\varphi^{\prime}(\xi)=\varphi_{1}\>,\end{cases}$ (1)
which yields highly oscillatory solutions. Here, $0<\varepsilon\ll 1$ is a very small parameter and $a$ is a real-valued function satisfying $a(x)\geq a_{0}>0$ and, for a quantum mechanical problem, it is related to the potential. The constants $\varphi_{0},\varphi_{1}\in\mathbb{C}$ may depend on $\varepsilon$ but are assumed to be ${\varepsilon}$-uniformly bounded. It is known that the (local) wavelength $\lambda$ of the solution $\varphi$ to (1) is proportional to $\varepsilon$. More precisely, it can be expressed as $\lambda(x)=(2\pi\varepsilon)/\sqrt{a(x)}$. Consequently, for a small parameter $\varepsilon$ the solution becomes highly oscillatory, particularly in the semi-classical limit $\varepsilon\to 0$.
Highly oscillatory problems such as (1) occur across a broad range of applications, e.g., plasma physics [13, 18], inflationary cosmology [20, 28] and electron transport in semiconductor devices such as resonant tunneling diodes [22, 26, 23]. More specifically, the state of an electron of mass $m$ that is injected with the prescribed energy $E$ from the right boundary into an electronic device (e.g., diode), modeled on the interval $[\xi,\eta]$, can be described by the following boundary value problem (BVP) (e.g., see [3] or [23, Chap. 2]):
$\begin{cases}-\varepsilon^{2}\psi_{E}^{\prime\prime}(x)+V(x)\psi_{E}(x)=E\psi_{E}(x)\>,\quad x\in(\xi,\eta)\>,\\ \psi_{E}^{\prime}(\xi)+\operatorname{i}k(\xi){\color[rgb]{0,0,0}\psi_{E}(\xi)}=0\>,\\ \psi_{E}^{\prime}(\eta)-\operatorname{i}k(\eta){\color[rgb]{0,0,0}\psi_{E}(\eta)}=-2\operatorname{i}k(\eta)\>.\end{cases}$ (2)
@@ 1.1 Background and approach
Since the solution $\varphi$ to (1) exhibits rapid oscillations when $\varepsilon$ is small, standard numerical methods for ODEs become inefficient as they are typically constrained by grid limitations $h=\mathcal{O}(\varepsilon)$ ($h$ denoting the step size), in order to resolve the oscillations accurately. By contrast, the phase function method of [8] is based on the observation that solutions to (1) can be represented accurately by means of a nonoscillatory phase function. Our approach presented below is closer to the uniformly accurate (w.r.t. $\varepsilon$) marching methods of [19, 16] which yield global errors of order $\mathcal{O}(h^{2})$ and allow to reduce the grid limitation to at least $h=\mathcal{O}(\sqrt{\varepsilon})$. The WKB-based (named after the physicists Wentzel, Kramers, Brillouin; cf. [17]) one-step method from [3] is even asymptotically correct, i.e. the numerical error goes to zero with $\varepsilon\to 0$, provided that the integrals $\int^{x}\sqrt{a(\tau)}\,\mathrm{d}\tau$ and $\int^{x}a(\tau)^{-1/4}(a(\tau)^{-1/4})^{\prime\prime}\,\mathrm{d}\tau$ for the phase of the solution can be computed exactly. More precisely, the method then yields an error which is of order $\mathcal{O}(\varepsilon^{3})$ as $\varepsilon\to 0$ and $\mathcal{O}(h^{2})$ as $h\to 0$. If these integrals cannot be evaluated exactly, the asymptotically correct error behavior can be (almost) recovered by employing spectral methods for the integrals, as shown in [4]. Further, in [2] the authors propose a numerical algorithm that switches adaptively between a defect correction iteration (which builds on an asymptotic expansion) for oscillatory regions of the solution, and a conventional Chebyshev collocation solver for smoother regions. Although the method is demonstrated to be highly accurate and efficient, a full error analysis was left for future work.
@@
The key question when implementing this WKB approximation is which choice of $N$ is adequate or even optimal, in the sense of minimizing the resulting approximation error. Indeed, since the asymptotic WKB series is typically divergent, the error cannot simply be reduced further by increasing the value of $N$. This question about the best attainable accuracy of the WKB approximation was already addressed in [28], where the author compared the WKB series with the exact solution represented by a convergent Bremmer series [9, 5], or more precisely, by an asymptotic expansion of that Bremmer series. The author finds that in cases where the coefficient function $a$ is analytic, the optimal truncation order is proportional to $\varepsilon^{-1}$, yielding a corresponding optimal accuracy which is exponentially small w.r.t. $\varepsilon$. However, to derive these results, the author makes several additional asymptotic approximations. In the present paper, on the other hand, we shall follow a more rigorous strategy by providing error estimates for the WKB approximation which are explicit w.r.t. $\varepsilon$ and $N$. We note, however, that the key assumption from [28], i.e., $a$ being analytic, will also be crucial for the strategy of the present work.
In practical finite precision computations, optimal truncation is not generally needed since it is not useful to add additional terms after reaching machine precision. In this paper we present concrete a-priori estimates for this truncation.
@@ 1.2 Main results
Our first main result is Theorem 3.7, which provides an explicit (w.r.t. $\varepsilon$ and $N$) error estimate for the WKB approximation, and implies that the approximation error is of order $\mathcal{O}(\varepsilon^{N})$. The explicitness of this estimate then allows the investigation of the error w.r.t. the truncation order $N$. Indeed, the optimal truncation order $N_{opt}$ can be predicted by minimizing the established upper error bound w.r.t. $N$ or by determining the smallest term of the asymptotic series, and is found to be proportional to $\varepsilon^{-1}$. This leads to our second main result, namely, Corollary 4.1. It states that, for an adequate choice of $N=N(\varepsilon)\sim\varepsilon^{-1}$, the error of the WKB approximation is of order $\mathcal{O}(\exp(-r/\varepsilon))$, $r>0$ being some constant. As a consequence, also the error of the optimally truncated WKB approximation is of order $\mathcal{O}(\exp(-r/\varepsilon))$, see also Remark 4.2.
@@ 2 WKB approximation
According to [6, Sec. 10.2], for the WKB-ansatz (5)-(6) to be valid on the whole interval $I$, it is necessary that the series $\sum_{n=0}^{\infty}\varepsilon^{n-1}S_{n}(x)$ is a uniform asymptotic expansion of $\varepsilon^{-1}S(x)$ as $\varepsilon\to 0$. This implies that for any $n\in\mathbb{N}_{0}$ the relation
@@ 3 Error analysis
with constants $\hat{\varphi}_{0},\hat{\varphi}_{1}\in\mathbb{C}$. Then there exists $C>0$ independent of $\varepsilon$ such that it holds for all $f\in C(I)$ and $\hat{\varphi}_{0},\hat{\varphi}_{1}\in\mathbb{C}$:
@@
$\frac{\mathrm{d}}{\mathrm{d}x}E(x)$ $\leq\frac{1}{\varepsilon^{2}}{\color[rgb]{0,0,0}|f(x)|^{2}}+cE(x)\>.$ (25)
Applying Gronwall’s inequality (i.e., multiply (25) by $\exp(-cx)$ and integrate), we therefore get
$E(x)$ ${\color[rgb]{0,0,0}\leq\frac{1}{\varepsilon^{2}}\int_{\xi}^{x}|f(s)|^{2}\operatorname{e}^{c(x-s)}\mathrm{d}s+E(\xi)\operatorname{e}^{c(x-\xi)}}$
$\leq\left(\frac{1}{\varepsilon^{2}}\lVert f\rVert_{L^{2}(I)}^{2}+E(\xi)\right)\operatorname{e}^{c(x-\xi)}$
@@
In order to derive an error estimate for the WKB approximation (13), which is explicit not only w.r.t. $\varepsilon$ but also w.r.t. the truncation order $N$, it is essential to control the growth of the functions $S_{n}$ w.r.t. $n\in\mathbb{N}_{0}$. As a first step, we aim to establish upper bounds for the derivatives $S_{n}^{\prime}$ which are given by recurrence relation (8)-(10). To this end, we employ a strategy similar to [21, Lemma 2], which relies heavily on Cauchy’s integral formula. To enable us to apply this tool, we shall assume that $S_{0}^{\prime}$ is not only defined on the real interval $I$, but also on a complex neighbourhood $G\subset\mathbb{C}$ of $I$ and holomorphic there. This leads us to introduce the following assumption.
@@
Let $S_{0}^{\prime}$ be holomorphic (complex analytic) on a complex, bounded, simply connected neighbourhood $G\subset\mathbb{C}$ of $I$, satisfying $S_{0}^{\prime}(z)\neq 0$ for any $z\in G$.
As a consequence of Hypothesis A, the function $a$ and all $S_{n}$, $n\in\mathbb{N}$, are holomorphic on $G$. In particular, each $S_{n}$ is bounded on $I$.
For the next lemma, we introduce, for $\delta>0$, the open sets
$G_{\delta}:=\{z\in G\mid\operatorname{dist}(z,\partial G)>\delta\}.$ (27)
@@
Let Hypothesis A be satisfied and let $0<\delta\leq 1$ be such that $G_{\delta}\neq\emptyset$. Then there exists a constant $K>0$ depending only on $G$ and $S_{0}^{\prime}$ such that
$\lVert S_{n}^{\prime}\rVert_{L^{\infty}(G_{\delta})}\leq\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}{\color[rgb]{0,0,0}K^{n}}n^{n}\delta^{-n}\>,\quad n\in\mathbb{N}_{0}\>.$ (28)
Here, we define $0^{0}$ as $1$.
@@
Note that since $S_{0}^{\prime}$ is holomorphic on $G$, it follows from recurrence relation (8)-(10) that $S_{n}^{\prime}$, and hence also $\widehat{S}_{n}^{\prime}$, is holomorphic on $G$, for every $n\in\mathbb{N}_{0}$. We will now prove by induction on $n$ that
$\lVert\widehat{S}_{n}^{\prime}\rVert_{L^{\infty}(G_{\delta})}\leq\frac{1}{2}{\color[rgb]{0,0,0}K^{n}}n^{n}\delta^{-n}\>,\quad n\in\mathbb{N}_{0}\>.$ (31)
Obviously, this estimate does hold for $n=0$, according to (29). Assume now that the estimate in (31) holds for $10\leq j\leq n-$ 1 with some fixed $n\geq 1$. We will now prove it for $n$. Let $10<\kappa<$ 1 and $z\in G_{\delta}$. We denote with $\partial B_{\kappa\delta}(z)$ a circle of radius $\kappa\delta$ around $z$, see the left panel of Figure 1.
@@
This, together with the fact that $\partial B_{\kappa\delta}(z)\subseteq{\color[rgb]{0,0,0}\overline{G_{(1-\kappa)\delta}}}$ yields
@@
$\leq\frac{1}{4}{\color[rgb]{0,0,0}K^{n}}\delta^{-n}\sum_{j=1}^{n-1}j^{j}(n-j)^{n-j}$
$\quad+\frac{1}{4}\lVert(S_{0}^{\prime})^{-2}\rVert_{L^{\infty}(G_{\delta})}\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}\delta^{-n}{\color[rgb]{0,0,0}K^{n-1}}\frac{(n-1)^{n-1}}{\kappa(1-\kappa)^{n-1}}\>.$ (34)
@@
$\lVert\widehat{S}_{n}^{\prime}\rVert_{L^{\infty}(G_{\delta})}$ $\leq\frac{1}{2}{\color[rgb]{0,0,0}K^{n}}n^{n}\delta^{-n}\left[\frac{1}{2}\left(\frac{n-1}{n}\right)^{n}+\frac{\lVert(S_{0}^{\prime})^{-2}\rVert_{L^{\infty}(G_{\delta})}\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}}{2{\color[rgb]{0,0,0}K}\kappa(1-\kappa)^{n-1}n}\left(\frac{n-1}{n}\right)^{n-1}\right]\>.$ (35)
@@
$\lVert\widehat{S}_{n}^{\prime}\rVert_{L^{\infty}(G_{\delta})}$ $\leq\frac{1}{2}{\color[rgb]{0,0,0}K^{n}}n^{n}\delta^{-n}\left[\frac{1}{2\operatorname{e}}+\frac{\lVert(S_{0}^{\prime})^{-2}\rVert_{L^{\infty}(G_{\delta})}\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}}{2{\color[rgb]{0,0,0}K}}\right]$ (36)
@@
${\color[rgb]{0,0,0}K:=\frac{\operatorname{e}}{2\operatorname{e}-1}\lVert(S_{0}^{\prime})^{-2}\rVert_{L^{\infty}(G)}}\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}\>.$ (37)
The estimate $\lVert(S_{0}^{\prime})^{-2}\rVert_{L^{\infty}(G_{\delta})}\leq\lVert(S_{0}^{\prime})^{-2}\rVert_{L^{\infty}(G)}$ concludes the proof. ∎
@@
By applying Lemma 3.2 on the r.h.s. of (40) we conclude that (39) holds with $K_{1}:=1/\delta$ and $K_{2}:={\color[rgb]{0,0,0}K/\delta}$. Estimate (38) then follows from (39) for $k=1$ and by the definition of $S_{n}$, see (14). ∎
@@
Of course, it is of great interest to find a constant $K_{2}$ from Corollary 3.3 is as small as possible. To this end one would have to minimize the constant $K/\delta$ in estimate (28). In particular, one has to fix some complex neighbourhood $G$ of $I$ as well as a constant $0<\delta\leq 1$ such that it holds $I\subset G_{\delta}$. Further, the proof of Lemma 3.2 indicates that $K$ can be reduced by choosing $G$ small, see (37). However, this means that one is forced to reduce also the value of $\delta$. Hence, this procedure usually results in a trade-off between the magnitudes of $K$ and $\delta$. More precisely, one is lead to solve the following minimization problem:
${\color[rgb]{0,0,0}\inf_{\begin{subarray}{c}0<\delta\leq 1\\ G\subset\mathbb{C}\\ I\subset G_{\delta}\end{subarray}}\frac{\lVert(S_{0}^{\prime})^{-2}\rVert_{L^{\infty}(G)}\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}}{\delta}=\min_{0<\delta\leq 1}\frac{\lVert(S_{0}^{\prime})^{-2}\rVert_{L^{\infty}(G^{\delta})}\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G^{\delta})}}{\delta}\>,}$ (41)
where $G^{\delta}:=\{z\in\mathbb{C}\mid\operatorname{dist}(z,I)<\delta\}$. Equality in (41) holds for the following reasons: First, on the l.h.s. of (41) one only needs to consider sets $G\subset\mathbb{C}$ of the form $G=G^{\kappa}:=\{z\in\mathbb{C}\mid\operatorname{dist}(z,I)<\kappa\}$, with $\kappa>0$. For any $0<\delta\leq 1$ such that $I\subset G_{\delta}$, this follows since the numerator on the l.h.s. of (41) is not increased when replacing $G$ by $G^{\operatorname{dist}(I,\partial G)}$, see the right panel of Figure 1. The condition $I\subset G_{\delta}$ then simply reads $\kappa>\delta$. Second, since for a fixed $0<\delta\leq 1$ and $\delta<\kappa_{1}<\kappa_{2}$ it holds that $G^{\kappa_{1}}\subset G^{\kappa_{2}}$, it is sufficient to consider simply the sets $G^{\delta+\epsilon}$, with $\epsilon>0$ being an arbitrarily small number. The equality in (41) then follows from the fact that ${\color[rgb]{0,0,0}\bigcap_{\epsilon>0}G^{\delta+\epsilon}=\overline{G^{\delta}}}$.
@@
$=\widetilde{\varphi}_{N}\left({\color[rgb]{0,0,0}\sum_{n=0}^{N}\sum_{k=0}^{N}\varepsilon^{n+k}S_{n}^{\prime}S_{k}^{\prime}}+\sum_{n=0}^{N}\varepsilon^{n+1}S_{n}^{\prime\prime}+a(x)\right)\>.$ (44)